OpenStudy (anonymous):

Find the equation of the line for the y-intercept and slope: (0,2), m=4

OpenStudy (anonymous):

use y-y1=m(x-x1) as the formula

OpenStudy (anonymous):

y-2=4(x-0) y-2=4x y=4x+2

OpenStudy (anonymous):

So if I need to mark on the graph a slope of 4, I go over 4 and down 1, right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Okay, thank you!

OpenStudy (anonymous):

NP

OpenStudy (anonymous):

so everytime they ask u to find the equation of a line, use that formula, that's if they give u the co-ordinates and the slope